a couple borrowed $450.00 from hometown Bank at 13.5% simple interest for 9 months for home repairs. to the nearest cent what is the total amount due after 9 months
Answer by stanbon(75887)

a couple borrowed $450.00 from hometown Bank at 13.5% simple interest for 9 months for home repairs. to the nearest cent what is the total amount due after 9 months
-------
A(t) = P(1+rt)
A(3/4) = 450(1+(0.135)*(3/4) = 450(1.1013) = $495.56
==================
Cheers,
Stan H.
==================